OpenStudy (anonymous):

F=1/7Pie m^2t solve for t

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok isthat the answer or part of showing the work?

OpenStudy (anonymous):

wow i made a huge mistake!

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

lets start again because what i wrote was way wrong. we want to make it look like \[b^x=A\] and then we solve via \[x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

\[F=\frac{1}{7}\pi m^{2t}\] \[7F=\pi m^{2t}\] \[\frac{7}{\pi}F=m^{2t}\] and so finally \[2t=\frac{\ln(\frac{\pi}{7}F)}{\ln(m)}\]

OpenStudy (anonymous):

divide both sides by 2 and you are done

OpenStudy (anonymous):

ok im sorry but i dont know how to divide on this can you show me?